the sum of two numbers is 62. their difference is 16. find the numbers.

Answer by Jstrasner(112)

Hey,
So for this we need to write down what we know and go from there:
x + y = 62
y = x -16
Now we simply substitute x - 16 for y in the first equation:
x + x -16 = 62 => 2x - 16 = 62 => 2x = 78 => x = 39
Now we substitute 39 for x in the second equation:
y = 39 - 16 => y = 23
So, x = 39, y = 23
I hope this helps!